the measure of the supplement of an angle is 50 less than 4 times the measure of the complement of the angle. Find the measure of the angle
Answer by cleomenius(959)

(180 - x) + 50 = 4(90-x)
180 - x + 50 = 360 - 4x
230 - x = 360 - 4x
3x = 130
x = 43.3 degrees, the angle.
complement = 46.7 degrees
supplement = 136.7 degrees
4(46.7)4 = 186.8; 186.8 - 50 = 136.7 degrees, check
Cleomenius.